# A Coordinated Routing Approach for Enhancing Bus Timeliness and Travel Efficiency in Mixed-Traffic Environment

### v1

@@ VI Conclusions
In this paper, we proposed a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment. A central element of our approach was a real-time traffic monitoring mechanism that estimated traffic flow and travel time on each road segment using sensor-derived data. Leveraging these estimates, we designed a coordinated scheme that identified a minimal subset of CAVs for rerouting and assigned them optimized paths reflecting current traffic conditions. By dynamically updating CAV routes in response to predicted congestion along dedicated lanes, our method effectively minimized disruptions to transit operations. Simulation experiments conducted in SUMO confirmed that the proposed strategy improved both transit reliability and CAV performance under diverse traffic scenarios. In future work, we plan to extend this framework to accommodate multi-modal transportation systems and assess their scalability in large-scale urban networks.

### v2

@@ VI Conclusions
In this paper, we presented a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment. A central element of our approach was a real-time traffic monitoring system that estimates traffic flow conditions on each road segment using sensor data. Leveraging these estimates, we developed a coordination scheme that identifies a minimal subset of CAVs for rerouting and assigns them optimized paths reflecting not only current but also anticipated potential congestion. By dynamically updating CAV routes in response to time-varying travel environments along DLs, our method effectively mitigates disruptions to transit operations. Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance under diverse traffic scenarios. Future work will focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban networks.